The rectangle shown is 3 times as long as it is wide. State algebraically that perimeter is 2 inches. What are the dimensions of the rectangle?
Answer by stanbon(75887)

The rectangle shown is 3 times as long as it is wide. State algebraically that perimeter is 2 inches. What are the dimensions of the rectangle?
-------------------------
Let width = x
Then length = 3x
-------
p = 2(length +width)
2 = 2(3x + x)
-------
1 = 4x
x = 1/4 inches (width)
3x = 3/4 inches (length)
===========================
Cheers,
Stan H.
=================